OpenStudy (anonymous):

A question of Limits. If f(x) approaches infinity as x approaches infinity, would you say that the limit of f(x) equals infinity or does not exist?

OpenStudy (anonymous):

You would notate it like so: \[\lim_{x\rightarrow\infty}{f(x)}=\infty\]

OpenStudy (anonymous):

I know how to notate it, what I am questioning is whether it is correct to have the "limit" of a function as infinity or if that limit wouldn't exist.

OpenStudy (anonymous):

Oh I see, well in that case it would not exist. Consider the following: \(f(x)=x\) \(g(x)=2x\) If you tried notating that it existed, you would have: \[\lim_{x\rightarrow\infty}{f(x)}=\infty\] And \[\lim_{x\rightarrow\infty}{g(x)}=\infty\] So you would be able to come to the conclusion that: \[\lim_{x\rightarrow\infty}{f(x)}=\lim_{x\rightarrow\infty}{g(x)}\] We know this to be incorrect since the two functions are in fact different so therefore, you would say that the limit does not exist